2018-02-06-MEMOMIPO » History » Version 4
Whyte, Colin, 06 February 2018 16:15
1 | 1 | Long, Kenneth | h1. 2018-02-06-MEMOMIPO |
---|---|---|---|
2 | |||
3 | h2. Joint MICE Experiment Management Office and MICE International Project Office |
||
4 | |||
5 | h3. 06th February 2018; 15:30 GMT |
||
6 | *RAL, R1 2nd Floor East, MICE Meeting room* |
||
7 | |||
8 | "Phone details":http://mice.iit.edu/phonebridge.html |
||
9 | |||
10 | # *Introduction: KL* |
||
11 | # *Notes of previous meeting and status of actions: All* |
||
12 | ** *CW:* Organise test of QD/QP systems for SSs. |
||
13 | 4 | Whyte, Colin | *** CMcW and JB to check this week. - no config files. |
14 | *** SSU at 3T and match coils - nominal settings |
||
15 | *** SSD 3T and then 2T. |
||
16 | *** Action CW - bracing bars. CW to discuss with JB. |
||
17 | 1 | Long, Kenneth | ** *KL:* Contact supervisor of CERN field-mapping team to seek to secure permission for them to carry out the MICE magnetic-field survey. |
18 | 4 | Whyte, Colin | *** complete agreed on week of 19th March |
19 | 1 | Long, Kenneth | ** *KL/CW:* Issue first version of the decommissioning plan this week. |
20 | 4 | Whyte, Colin | *** completed today. |
21 | 1 | Long, Kenneth | # *Hall status, stability of systems, planning for decommissioning and making safe to be complete by 31Mar18: CW* |
22 | 4 | Whyte, Colin | *** Move out R9 to Hall |
23 | *** Quads, Tof2 KL, EMR. |
||
24 | *** Action CW to contact Oxford Machine shop re-diffuser. |
||
25 | *** FC pressure - helium pressure 1160mbar - cryo mech 340/310psi - high ambient pressure. Leave until Mon 11th. |
||
26 | *** SSD pressure increase next week. |
||
27 | *** US DoE |
||
28 | **** Mu2E QPS system |
||
29 | **** cryocoolers and compressors - Mu2E - request is from institution - not approved at higher level. |
||
30 | **** MP to raise with DoE |
||
31 | 1 | Long, Kenneth | # *Data validation ahead of start of Cycle 2017/04: CR* |
32 | 3 | Rogers, Chris | ** [[analysis:2018-02-01_no-run_summary|Summary of studies of data validity]] |
33 | 1 | Long, Kenneth | # *Software and computing status: DR* |
34 | 4 | Whyte, Colin | ** issues on processing time - run on RAL PPD grid. Making changes with Chris Brew - should be on line by next week. |
35 | ** memory issues solved |
||
36 | ** some geometry issues - DR working to solve. Most recent survey was before wedge change. |
||
37 | ** tracker and ToF meetings still to be scheduled |
||
38 | ** computing infrastructure inventory in progress. |
||
39 | 1 | Long, Kenneth | # *Memo on measurements required before experiment is decommissioned: CR* |
40 | 2 | Rogers, Chris | ** No update |
41 | 1 | Long, Kenneth | # *Paper status and plans: CR* |
42 | 4 | Whyte, Colin | ** meetings last week |
43 | *** scattering straight tracks steady progress - refs meeting at CM |
||
44 | *** VB no progress report - ready for CM |
||
45 | *** CR plan to be ready for CM50 |
||
46 | 1 | Long, Kenneth | # *AOB* |
47 | |||
48 | h4. Summary of actions: |
||
49 | |||
50 | |||
51 | h3. Back to MEMO: |
||
52 | |||
53 | [[memo:|Back]] |
||
54 | |||
55 | h3. Back to MIPO: |
||
56 | |||
57 | [[mipo:|Back]] |
||
58 | |||
59 | h2. Documents |